### v1

# Primordial Black Holes from Primordial Voids

Abstract
Primordial black holes (PBHs) are a compelling dark matter candidate and a unique probe of small-scale cosmological fluctuations. Their formation is usually attributed to large positive curvature perturbations, which collapse upon Hubble re-entry during radiation domination. In this work we investigate instead the role of negative curvature perturbations, corresponding to the growth of primordial void (PV) like regions. Using numerical relativity simulations, we show that sufficiently deep PV can undergo a nonlinear rebounce at the center, generating an effective overdensity that eventually collapses into a PBH. We determine the critical threshold for this process for a variety of equations of state, and demonstrate that the resulting black holes obey a scaling relation analogous to the standard overdensity case. These results establish primordial voids as a novel channel for PBH formation and highlight their potential impact on PBH abundances and cosmological signatures.
pacs:
98.80.Cq, 98.70.Vc
I Introduction
Primordial black holes (PBHs) Hawking (1971); Carr and Hawking (1974) are among the most promising candidates to explain the nature of dark matter, while also providing a unique probe of physics in the early Universe. Unlike astrophysical black holes, PBHs arise from non-stellar processes operating shortly after inflation, from processes that include the collapse of large adabiatic and iso-curvature perturbations  Shibata and Sasaki (1999); Passaglia and Sasaki (2022); Yoo et al. (2022), cosmic string loops Jenkins and Sakellariadou (2020); Helfer et al. (2019), or bubble collisions Jung and Okui (2024) (for reviews see Bagui et al. (2025); Escrivà et al. (2022); Carr and Kuhnel (2020)). Among all possibilities, the most widely studied and theoretically well-motivated scenario involves the collapse of curvature fluctuations during the radiation-dominated era Polnarev and Musco (2007); Escrivà et al. (2020, 2021), and other pressure(less) fluids Mack et al. (2007); Harada et al. (2016). These perturbations are typically seeded by quantum fluctuations during inflation, stretched to super-Hubble scales, which later re-enter the Hubble volume.
In this standard picture, attention has centered on positive curvature perturbations, corresponding to overdense regions of spacetime. For such profiles, the local density contrast grows non-linearly upon Hubble re-entry, and collapse occurs if a critical threshold is exceeded Carr (1975); Musco et al. (2005); Musco and Miller (2013); Escrivà et al. (2020, 2021). This framework underlies the bulk of PBH formation models, and besides being natural dark matter candidates, they can play an important role as seeds of supermassive black holes and be the sources of gravitational waves detected by the LIGO-Virgo-KAGRA experiments Sasaki et al. (2018); Carr and Kuhnel (2020); Bagui et al. (2025).
Figure 1: Amplitude thresholds for type-A PBH formation across different equations of state $\omega$. Black dots denote results from numerical relativity simulations, connected by a solid black line (quadratic interpolation) representing the separatrix. The light and dark red (green) shaded regions correspond to type-I and type-II classifications for positive (negative) curvature fluctuations, respectively.
By contrast, the role of negative curvature perturbations has received little attention. These fluctuations correspond to underdensities or void-like regions in the primordial plasma, which we refer to as Primoridal Voids (PVs). At first glance, they appear unlikely to collapse, since their geometry counteracts gravitational attraction. However, once the nonlinear dynamics of Hubble re-entry are taken into account, deep negative curvature perturbations may evolve in a qualitatively different way. Rather than simply dispersing, such void-like profiles can generate overdense shell-structures that, in turn, might drive gravitational collapse and produce PBHs. This provides a new and unexplored channel for PBH formation, complementing the standard overdensity case.
The study of this mechanism requires going beyond linear perturbation theory. The general relativistic evolution of such void-like perturbations involves strong nonlinearities both in the metric and in the fluid sector, which cannot be captured analytically in a closed form. Numerical relativity offers the necessary framework to follow these dynamics from Hubble entry through to possible collapse, while analytical arguments can guide the identification of thresholds and scaling laws. Together, these tools enable a systematic exploration of PBH formation from negative curvature perturbations.
In this work we carry out for the first time the studies on PBH formations from rebounces of PVs. We identify the critical amplitude of negative curvature fluctuations required for collapse, demonstrate that the resulting PBHs obey a scaling relation akin to the overdensity case, and map the competing outcomes of dispersal, shock formation, and black-hole collapse. These results open a new avenue for PBH phenomenology and highlight the need to include PVs when assessing PBH abundances and their observational consequences.
The paper is organized as follows. In Section II, we review the theoretical framework of PBH formation from curvature fluctuations. Section III details our numerical methodology based on general relativistic hydrodynamics. Our main results on collapse dynamics, critical thresholds, and mass scaling are presented in Section IV. The conclusions are left for Section V. Throughout this work, we use the “mostly-plus” metric signature (,+,+,+)(-,+,+,+) and assume geometrical units where $G=c=1$.
II Primordial black holes from curvature fluctuations
Large curvature perturbations in the early Universe can arise from inflationary dynamics beyond the simplest slow-roll picture. In particular, non-attractor phases provide a natural way to amplify the curvature perturbation $\zeta$ on super-Hubble scales. Examples include ultra-slow roll inflation Figueroa et al. (2021); Pattison et al. (2021), constant-roll inflation Motohashi et al. (2020); Motohashi and Starobinsky (2017a, b); Tomberg (2023); Inui et al. (2024), or spectator field scenarios such as curvaton dynamics Pi and Sasaki (2023a). In these cases, the usual conservation of $\zeta$ outside the Hubble volume is broken, allowing its amplitude to grow and potentially reach the large values required for PBH formation. A useful organizing principle for these scenarios is provided by the so-called logarithmic duality Pi and Sasaki (2023b); Inui et al. (2025, 2024), which relates the non-Gaussian curvature perturbation to an underlying Gaussian field and connects various non-attractor mechanisms. Explicitly, the curvature takes the form
$\zeta_{\rm log}(\vec{x})\;=\;-\frac{1}{\gamma_{\rm NG}}\,\log\!\left(1-\gamma_{\rm NG}\,\zeta_{G}(\vec{x})\right)\,$, (1)
where $\zeta_{G}$ denotes the Gaussian curvature perturbation and $\gamma_{\rm NG}$ parametrizes the strength and sign of non-Gaussian effects. Besides modifying the mean curvature profile, this relation also alters the probability distribution function. Models with $\gamma_{\rm NG}>0$ introduce a positive skewness that enhances the abundance of overdensities, whereas models with $\gamma_{\rm NG}<0$ enhance the population of voids.
For simplicity, however, we adopt a set of assumptions that notably reduce the complexity of the problem. First, we restrict ourselves to the Gaussian case ($\gamma_{\rm NG}=0$), ignoring non-Gaussian modification to the shape profile. Second, we assume a narrow-peak (monochromatic) power spectrum, with amplitude $A_{\zeta}$ peaked at $k_{\star}$,
$\mathcal{P}_{\zeta}(k)=A_{\zeta}\delta(k-k_{\star})~,$ (2)
which reduces scale dependence of the system to $k_{\star}$. Finally, we focus on the rare large peaks (maxima or minima) where the spherical symmetry approximation is justified. Under these assumptions, the mean radial profile of a curvature perturbation is given by a sinc-function Bardeen et al. (1986); Yoo et al. (2018, 2021),
$\zeta(r)\;=\;\mu\,\mathrm{sinc}\!\left(k_{\star}r\right)\,$, (3)
where $\mu$ is the central amplitude. The study of other shape profiles—such as those linked to a broad-peak power spectrum or the effects of non-Gaussianities—is beyond the scope of this paper and is left for future work. Here, we focus on the profile given by Eq. (3) and use it to generate the initial conditions for the numerical relativity simulations presented in Section IV.
II.1 Geometric classifications of fluctuations and dynamical collapse
Figure 2: Initial configuration for three representative negative curvature fluctuations with amplitudes $\mu=-0.6$ (red), 1-1 (black), and 1.6-1.6 (green). Top panels: the initial curvature profile $\zeta(r)$, the areal radius R(r){R}(r), and the three-Ricci scalar $^{(3)}\,R(r)$. In the latter panel, positive values are represented with solid lines, whereas negative values with dashed line. Bottom panels: the corresponding non-linear compaction function $\mathcal{C}(r)$, its linear component, and the compaction’s curvature $\kappa(r)$, computed assuming radiation domination ($\omega=1/3$). The chosen amplitudes exemplify the type-I (red), type-I/II transitional (black), and type-II (green) regimes. The vertical blue line in the bottom panels indicate the location of the maximum of $C_{\ell}$.
An important quantity for characterizing the geometry of a curvature fluctuation is the areal radius, defined as
$R(r)=\mathfrak{a}(t)e^{\zeta(r)}r,$ (4)
where $r$ is the comoving radius and $\mathfrak{a}(t)$ is the cosmological scale factor. Using the notation from Refs. Kopp et al. (2011); Uehara et al. (2024); Harada et al. (2025), we can classify super-Hubble curvature fluctuations as either type I or type II. In type I, $R(r)$ is monotonic, such that $\mathrm{d}R/\mathrm{d}r>0$ everywhere, whereas in the type II case, the areal radius develops a non-monotonic region where $\mathrm{d}R/\mathrm{d}r<0$. This non-monotonicity leads to a warped geometry, developing a ”belly” and ”neck” structure, where the angular distance between diametrically opposite points inside the ”neck” becomes shorter than the radial distance between them. For the sinc profile with $\mu>0$, the transition between type I and type II occurs at $\mu\gtrsim 0.95$. Notably, negative amplitude fluctuations also admit type II configurations, which occur for $\mu<-1$ (see Fig. 2 for representative examples).
It is important to stress that the classification into type I and type II refers to the geometry of the super-Hubble fluctuation, while the recently introduced classification into type-A and type-B instead concerns a dynamical classification of the collapse process of a PBH. Type-A PBHs form after Hubble re-entry, when the central overdensity grows sufficiently to overcome pressure support and collapse, eventually creating an apparent horizon. Type-B PBHs, in contrast, form through a ”separate universe” mechanism. Here, a sufficiently strong curvature fluctuation creates a causally disconnected region from the background expansion, which develops a trapped surface before being incorporated into the Hubble volume of the expanding background (see Shimada et al. (2024); Uehara et al. (2024) for recent studies). In this sense, type A/B is a dynamical classification of the collapse process, whereas type I/II is a geometric classification of the curvature profile at super-Hubble scales.
As recently proposed in  Germani and Montellà (2025), an additional relevant classification concerns the sign of the spatial curvature at the fluctuation’s core, providing a complementary geometric characterization than the Type I/II classification. We can use the three-dimensional Ricci scalar to classify the fluctuation’s core geometry as open, flat, or closed. At super-Hubble scales, the three-Ricci scalar reads
$^{(3)}\,\mathrm{R}=-e^{-2\zeta(r)}\left(4\zeta^{\prime\prime}(r)+\frac{8}{r}\zeta^{\prime}(r)+2\left(\zeta^{\prime}(r)\right)^{2}\right).$ (5)
This classification is relevant for PBH formation, because a closed core would facilitate the collapse, leading to a lower critical amplitude threshold, whereas an open core would oppose it. In this light, positive and negative values of $\mu$ in Eq.(3) correspond to open and closed core-types, respectively, and therefore a higher collapse threshold is expected for the PVs sector.
II.2 Trapped Surfaces and Misner–Sharp Mass in an Expanding Universe
The local geometry of a spacetime hypersurface can be characterized by the null expansions of the outgoing and ingoing radial null congruences, $\Theta_{+}$ and $\Theta_{-}$, respectively.
For instance, in the pure expanding FLRW universe, the null expansions are defined as
$\Theta_{\pm}=\frac{2}{R}\left(HR\pm 1\right)~,$ (6)
and therefore, the structure of null congruences reads
$\Theta_{+}>0,\ \Theta_{-}<0\quad$ $\text{for}\quad HR<1,$ (7)
$\Theta_{+}>0,\ \Theta_{-}=0\quad$ $\text{for}\quad HR=1,$ (8)
$\Theta_{+}>0,\ \Theta_{-}>0\quad$ $\text{for}\quad HR>1.$ (9)
The first case corresponds to an untrapped region. The second case, at $\Theta_{-}=0$, defines the marginally outer trapped surface, which coincides with the Hubble radius. For $HR>1$, the $\Theta_{-}$ become positive, corresponding to a past-trapped region.
A covariant characterization of the gravitational energy contained within a volume is provided by the Misner–Sharp mass Misner and Sharp (1964); May and White (1966), defined as
$M_{\mathrm{MS}}=\frac{R}{2}\left(1-g^{\mu\nu}\partial_{\mu}R\,\partial_{\nu}R\right).$ (10)
The Misner–Sharp mass satisfies the geometric relation
$1-\frac{2M_{\mathrm{MS}}}{R}=g^{\mu\nu}\partial_{\mu}R\,\partial_{\nu}R,$ (11)
which vanishes at marginally trapped surfaces. Thus, the condition
$\frac{2M_{\mathrm{MS}}}{R}=1$ (12)
defines the existence of an apparent horizon (AH), corresponding to the null expansion condition $\Theta_{+}\,\Theta_{-}=0$. Thus, the condition (12) is satisfied both at the Hubble distance of the cosmological background for the $\Theta_{-}=0$ case (a marginally past-trapped surface), and at the local AH of a forming PBH corresponding to $\Theta_{+}=0$ (a marginally future-trapped surface).
After a PBH is formed, the causal structure of null congruences changes towards to
$\Theta_{+}<0,\ \Theta_{-}<0\quad$ $\text{for}\quad R<R_{\text{PBH}},$ (13)
$\Theta_{+}=0,\ \Theta_{-}<0\quad$ $\text{for}\quad R=R_{\text{PBH}},$ (14)
$\Theta_{+}>0,\ \Theta_{-}<0\quad$ $\text{for}\quad R_{\text{PBH}}<R<R_{H},$ (15)
$\Theta_{+}>0,\ \Theta_{-}=0\quad$ $\text{for}\quad R=R_{H},$ (16)
$\Theta_{+}>0,\ \Theta_{-}>0\quad$ $\text{for}\quad R>R_{H},$ (17)
where $R_{\text{PBH}}$ denotes the AH of the PBH, and $R_{H}$ the Hubble radius.
II.3 Compaction function and threshold criterion
An alternative diagnostic for identifying the formation of a black hole is the compaction function, which can be expressed as
$C(r)=\frac{2[M_{\rm MS}(r)-M_{\rm bkg}(r)]}{R(r)},$ (18)
where $M_{\rm bkg}(r)$ is the FLRW background mass within the same radius. When $C(r)\gtrsim 1$, a trapped surface forms, signaling black hole formation.
An important application of the compaction function is that its peak value at superHubble scales can be used to predict whether positive amplitude fluctuation will collapse into a black hole upon Hubble re-entry. Following Refs. Germani and Sheth (2023), one can define the linear compaction function as
$C_{\ell}(r)=2\phi(w)\,r\,\zeta^{\prime}(r)\,,\quad\phi(w)\equiv\frac{(1+3w)}{5+3w}\,$, (19)
which depends only on the profile of $\zeta(r)$ and the equation-of-state parameter $w$. The full (non-linear) compaction function is then given by
$C(r)=C_{\ell}\left(1-\frac{1}{4\phi(w)}C_{\ell}\right).$ (20)
It is useful to introduce the curvature of the compaction function,
$\kappa(r)\equiv-r^{2}\frac{\mathrm{d}^{2}C_{\ell}}{\mathrm{d}r^{2}}~,$ (21)
and to evaluate this quantity at the radius $r_{m}$ where $C_{\ell}(r)$ attains its maximum. Let us use the notation $C_{m}\equiv C_{\ell}(r_{m})$ and $\kappa_{m}\equiv\kappa(r_{m})$ for brevity.
It has been shown in Escrivà et al. (2021, 2020) that the critical threshold for primordial black hole formation follows from a non-linear relation between $C_{m}$ and $\kappa_{m}$. This yields an implicit criterion of the form
$C_{m}>\delta_{c}(C_{m},\kappa_{m}),$ (22)
where $\delta_{c}(C_{m},\kappa_{m})$ is a known non-linear relation (see Escrivà (2025) for the type II case). This approach has been shown to reproduce the thresholds measured in full numerical-relativity simulations with high accuracy (within the 5%) for the radiation domination case, i.e. $\omega=1/3$.
III General Relativistic Hydrodynamics
Figure 3: Evolution of the density contrast $\rho_{E}/\rho_{\rm bkg})$ for negative curvature fluctuations re-entering the Hubble horizon during radiation domination $(\omega=1/3)$, obtained from numerical relativity simulations. Top panels: initial fluctuation with amplitude $\mu=-1.8$, which forms a PV that rebounds into a central overdensity and collapses into a PBH. Bottom panels: initial fluctuation with amplitude $\mu=-0.8$, which also forms a PV but whose rebound dissipates into sound waves. Underdensities and overdensities are represented by a blue-to-red colormap. The green circle denotes the formation of an apparent horizon.
We now describe the general relativistic hydrodynamics methods used to evolve Hubble re-entry of negative curvature fluctuations, the formation of PVs and potential collapse into PBHs.
We describe the cosmological system using the Baumgarte-Shapiro-Shibata-Nakamura (BSSN) formalism Baumgarte and Shapiro (1998); Shibata and Nakamura (1995), reduced to spherical symmetry Alcubierre and Mendez (2011). Thus, the line element reads
$ds^{2}=-\alpha^{2}dt^{2}+e^{4\chi}\left[a(dr+\beta^{r}dt)^{2}+r^{2}b\,d\Omega^{2}\right],$ (23)
where $\Omega$ is the solid angle, $\alpha$ is the lapse function, $\beta^{r}$ is the radial component of the shift vector. The spatial metric is conformally decomposed $\gamma_{ij}=e^{4\chi}\tilde{\gamma}_{ij}$, where $\chi$ is the conformal factor and the metric components are defined by $a(r,t)\equiv\tilde{\gamma}_{rr}$ and $b(r,t)\equiv\tilde{\gamma}_{\theta\theta}/r^{2}$.
The extrinsic curvature $K_{ij}$ is decomposed into its trace, $K$, and its traceless part, $\tilde{A}_{ij}$. In this formulation, the mixed components $A_{a}\equiv\tilde{A}^{r}_{r}$ and $A_{b}\equiv\tilde{A}^{\theta}_{\theta}$ are used, which are related by the traceless condition $A_{a}+2A_{b}=0$. The BSSN formalism also introduces the contracted Christoffel symbols of the conformal metric, $\hat{\Delta}^{i}\equiv-\partial_{j}\tilde{\gamma}^{ij}$, as independent variables, used for numerical stability in the simulations.
We choose the matter sector to be described by a perfect fluid with stress-energy tensor
$T_{\mu\nu}=(\rho_{\rm fl}+p)u_{\mu}u_{\nu}+pg_{\mu\nu},$ (24)
where $\rho_{\rm fl}$ and $p$ are the fluid’s energy density and pressure, respectively, and $u_{\mu}$ is the four-velocity. We employ a barotropic equation of state $p=\omega\rho_{\rm fl}$, considering values of $\omega$ spanning from $0<\omega<1$, including the radiation domination case with $\omega=1/3$. For a comoving observer, the energy density splits into rest energy density $\rho_{0}$ and internal energy $\varepsilon$ as $\rho_{\rm fl}\equiv\rho_{0}(1+\varepsilon)$. The energy density measured by an Eulerian observer is given by
$\rho_{\rm{E}}=(\rho_{\rm fl}+p)W^{2}-p,$ (25)
where $W\equiv 1/\sqrt{1-v^{2}}$ is the Lorentz factor and $v^{r}$ is the radial component of the fluid’s three-velocity.
The state of the fluid is described by the primitive variables: the rest mass density $\rho_{0}$, specific internal energy $\varepsilon$, pressure $p$, and three-velocity $v^{r}$. For numerical evolution, these are transformed into the conserved variables $D$, $S_{r}$, and $\mathcal{E}$, which are defined as
$D$ $=\rho_{0}W,$ (26)
$S_{r}$ $=(\rho_{\rm fl}+p)W^{2}v_{r},$ (27)
$\mathcal{E}$ $=(\rho_{\rm fl}+p)W^{2}-p-D~.$ (28)
The system is evolved by coupling the BSSN equations for the gravitational variables $\{\chi,a,b,K,A_{a},\hat{\Delta}^{r},\alpha,\beta^{r}\}$ with the hydrodynamic equations for the conserved matter variables $\{D,S_{r},\mathcal{E}\}$. The complete set of evolution equations is detailed in Ref. Alcubierre (2008); Alcubierre and Mendez (2011) and summarized in the Appendix. All simulations begin from constraint-satisfying initial conditions. The methods for solving the initial data, evolution schemes and other numerical details can be found in the Appendix.
IV Numerical results
Our numerical simulations begins from a initial equal-time hypersurface representing a super-Hubble curvature fluctuation $\zeta(t,r)$ centered at the origin of our radial coordinate, embedded in a Friedmann–Lemaître–Robertson–Walker (FLRW) expanding background. The overall scale factor is initially set to one, $\mathfrak{a}(t_{0})=1$ with an expanding Hubble rate $H(t)=\frac{2}{3(1+\omega)}\frac{1}{t}$, where $t_{0}$ is an arbitrary initial time in code units. The characteristic scale of $\zeta(t_{0},r)$ is set to $k_{\star}^{-1}=100\,H^{-1}(t_{0})$, ensuring that the perturbation is well outside the Hubble horizon initially.
IV.1 Dynamics of gravitational collapse
Figure 4: Mass scaling of PBH formation for negative (left pannel) and positive (right pannel) curvature fluctuations.
For completeness, we present results from both positive and negative $\zeta$ perturbations and compare them.
For a positive amplitude $\mu$, the perturbation corresponds to an initial overdensity. While outside the Hubble horizon, the perturbation remains quasi-linear. Its subsequent evolution after horizon entry is characterized by rapid growth and strong self-gravity. The final outcome is then determined by the competition between this self-gravity and pressure gradients. For sufficiently large amplitude $\mu>\mu_{c}^{(+)}$, the gravity dominates and the region collapses to form a PBH, whereas when not, the pressure support is sufficient to prevent the collapse and the overdensity disperses into sound waves without forming a black hole.
In contrast, a negative amplitude $\mu$ corresponds to an initial underdensity. The central underdensity deepens progressively, reaching its maximum depth at Hubble crossing—a stage we refer to as a PV. Owing to the oscillatory structure of the sinc profile, a PV is always accompanied by a surrounding overdense shell. During the subsequent sub-Hubble evolution, this shell contracts under its self-gravity and the background pressure, driving matter inwards and causing the central region to rebound. In this scenario, pressure initially drives the contraction, in direct contrast to the positive $\mu$ case where it always opposes collapse. For a sufficiently negative amplitude $\mu<\mu_{c}^{(-)}$, the rebound is sufficiently strong to collapse into a PBH; otherwise, it disperses, generating sound waves. This mechanism highlights the qualitative difference between PBH formation from overdensities and PVs. Figure 3 illustrates the dynamics of PVs in both collapsing and dissipating scenarios, showing the energy density ratio $\rho_{E}/\rho_{\rm bkg}$ for times before Hubble crossing, at rebound, and during the final collapse or dissipation of the central feature.
IV.2 Equation of state dependence and thresholds
The formation dynamics exhibit a strong dependence on the fluid pressure, parameterized by the equation of state $\omega$. We have computed the critical thresholds for PBH formation (see Table 1) from both overdensities and PVs across a range of $\omega$, from the nearly pressureless limit ($\omega\simeq 0$) to the stiff fluid limit ($\omega=1$). Figure 1 shows the numerical collapse thresholds $\mu_{\rm c}$ for representative $\omega$ values. As expected, the thresholds for PVs exceed those for overdensities, as gravitational collapse must overcome the expanding void geometry. Furthermore, we find that while positive-$\mu$ perturbations can form PBHs directly from Type I fluctuations, negative-$\mu$ perturbations require Type II fluctuations. All PBHs produced in the vicinity after $\left|\mu\right|>\left|\mu_{c}^{(\pm)}\right|$ are of Type A.
Furthermore, the thresholds for positive $\mu$ increase monotonically with $\omega$, elucidating the role of pressure to prevent gravitational collapse. Instead, thresholds in the PV formation channel display a non-monotonic behavior, peaking near the radiation-dominated case $\omega\simeq 1/3$ and decreasing for both softer and stiffer equations of state. We interpret this as a reflection of competing effects of pressure gradients and sound speed in the collapse dynamics. For $\omega<1/3$, the reduced fluid’s pressure facilitates the collapse once the rebounce is produced. Conversely, for $\omega>1/3$, the increased sound speed associated with higher pressure enables a faster compression rate, producing a more violent rebound that effectively facilitates PBH formation, even though pressure subsequently opposes collapse once the overdense core is generated.
Another example to illustrate the sound speed effect is on the near-dust limit ($\omega\to 0$). In this case, the overdense shell surrounding a PV barely contracts, but instead, the shell begins collapsing at nearly constant radius (i.e. the shell narrows in thickness rather than undergoing significant inward motion). Numerical limitations prevent definitive conclusions in this regime, but upon inspection, we see that the compaction function remains small at all times, suggesting that a PBH might never form. The fate of the thin shell formation would be presumably strongly conditioned by (aspherical) environmental effects which are not captured in our simulations.
After comparison of the numerical thresholds with those predicted by the analytical prescription (22), our numerical results confirm a good agreement for the positive-curvature fluctuations scenario in the radiation domination case, with deviations under 5%. For negative-curvature fluctuations, however, we observe that the presciption predicts a lower threshold ($\mu_{c}^{\rm th}\sim-1.1$), inducing approximately a 20% error. This disagreement is not surprising as these analytical approaches do not capture the specific open-core geometry associated to PVs.
IV.3 Scaling of mass functions
Finally, we compute the critical mass scaling for $\omega=1/5,1/3,3/5$, corresponding to soft, radiation, and stiff equations of state, respectively. The numerical results are fitted to the standard critical scaling form Choptuik (1993); Evans and Coleman (1994); Musco and Miller (2013),
$\frac{M_{\rm PBH}}{M_{\times}}=K\left|\mu-\mu_{c}\right|^{\gamma}~,$ (29)
where $M_{\times}$ is the mass within the Hubble volume at the time of Hubble-crossing, $K$ is a fitted mass ratio, $\gamma$ is the critical exponent, and $\mu_{c}$ is the critical threshold. The best fit values are shown in Table 2. Notably, we confirm a very similar fitting values between positive and negative curvature fluctuations at a given $\omega$, particularly for the scaling index $\gamma$ (see also Figure 4). The fits are performed, as customary, for near-threshold amplitudes $\delta\mu\equiv\left|\mu-\mu_{c}\right|<10^{-1}$, successfully recovering the well-known value of $\gamma\approx 0.36$ for radiation domination Evans and Coleman (1994). We observe, however, that for larger amplitudes $\delta\mu>0.1$, the resulting masses exceed the predictions of the critical scaling law, a deviation already seen in Escrivà et al. (2022).
Table 1: Numerical critical thresholds for several equations of state. Precision is $\pm 0.01$.
$\omega$ $\mu_{c}^{(+)}$ $\mu_{c}^{(-)}$
0.12 0.36 -1.10
0.20 0.47 -1.41
$1/3$ 0.60 -1.59
0.40 0.65 -1.60
0.60 0.70 -1.47
0.80 0.71 -1.30
1.00 0.72 -1.15
Table 2: Fitted parameters for different equations of state $\omega$ for $\mu>0$ and $\mu<0$.
   $\omega$ Fit parameters for $\mu>0$
   $3/5$ $K=9.18,\ \gamma=0.56,\ \mu_{c}^{(+)}=0.700$
   $1/3$ $K=8.44,\ \gamma=0.36,\ \mu_{c}^{(+)}=0.604$
   $1/5$ $K=7.87,\ \gamma=0.24,\ \mu_{c}^{(+)}=0.477$
   $\omega$ Fit parameters for $\mu<0$
   $3/5$ $K=9.04,\ \ \gamma=0.54,\ \mu_{c}^{(-)}=-1.469$
   $1/3$ $K=12.99,\ \gamma=0.35,\ \mu_{c}^{(-)}=-1.588$
   $1/5$ $K=16.96,\ \gamma=0.22,\ \mu_{c}^{(-)}=-1.409$
V Conclusions and Discussions
In this work we have investigated PBH formation from large negative curvature perturbations, what we dubbed as Primordial Voids. Using numerical relativity simulations, we demonstrated that sufficiently deep PVs undergo a nonlinear rebound at their center that triggers black hole formation. We report the amplitude thresholds and PBH mass scaling for several equations of state. The thresholds lie in the type-II region fluctuations, while the PBH collapse formed in the vicinity after these thresholds correspond to type-A. For a comparative reference, we have also extended our investigation to the case of positive-amplitude fluctuations.
This work represents one of the first investigations into the previously overlooked phenomenology of underdensities that may have originated during cosmic inflation. Several interesting extensions remain for future work, including studying the effects of non-Gaussianities on the shape profile—particularly those that produce a negative skewness in the probability distribution function of $\zeta$, as these would enhance the abundance of PVs and their imprints in cosmological observations.
In addition, the question of whether PVs can form PBHs of type-B (from a ”separate universe”) has been left unanswered. Would these objects exist, they would correspond (at least initially) to a trapped region with an open-core geometry. It would be interesting to better undertenant the fate of such structures.
Because underdensities or voids are necessarily present in any model that sufficiently enhances the scalar powerspectrum, it is important to take into account their associated phenomenology. Some examples include the generation of sound waves and related gravitational wave signatures Ning et al. (2025); Zeng et al. (2025), the potentially relevant effects to matter-dominated epochs such for PBH reheating models Allahverdi et al. (2025), or even their potential role as catalysts for baryogenesis Baumann et al. (2007); Carr et al. (2021); García-Bellido et al. (2021); Despontin et al. (2024).
Finally, extending this analysis to scalar-field-dominated cosmologies would also be relevant for scenarios where PBHs form shortly after inflation, either during a phase of kinationCheng et al. (2025), or (p)reheating Khlopov et al. (1985); Martin et al. (2020); Padilla et al. (2025); Milligan et al. (2025).
Acknowledgments
The authors thank Cristiano Germani, Shi Pi, Diego Cruces, Sébastien Clesse, Jaume Garriga, Shao-Jiang Wang, Xiang-Xi Zeng and Zhuan Ning for stimulating converstations. C. J. is supported by NSFC grants No. W2433007 and 12475066. Z.-Y. Y. is supported by an appointment to the Young Scientist Training (YST) program at the APCTP through the Science and Technology Promotion Fund and Lottery Fund of the Korean Government. This was also supported by the Korean Local Governments-Gyeongsangbuk-do Province and Pohang City. Computational resources were borrowed from the SCNet High-Performance Computing Infrastructure, Sugon (Dawning Information Industry Co., Ltd.)
Appendix A Details of the Numerical Relativity Simulations
A.1 Evolution equations
We solve the Einstein Field equations using the BSSN formalism, as described in Section III.
The evolution equations of the geometrical BSSN variables $\{\chi,a,b,K,A_{a},\hat{\Delta}^{r},\alpha,\beta^{r}\}$ are the following Alcubierre and Mendez (2011),
$\partial_{t}\chi$ $=$ $\beta^{r}\partial_{r}\chi-\frac{1}{6}\alpha K\;,$ (30)
$\partial_{t}a$ $=$ $\beta^{r}\partial_{r}a+2a\partial_{r}\beta^{r}-2\alpha aA_{a},\qquad$ (31)
$\partial_{t}b$ $=$ $\beta^{r}\partial_{r}b+2b\>\frac{\beta^{r}}{r}-2\alpha bA_{b}\;.$ (32)
$\partial_{t}K$ $=$ $\beta^{r}\partial_{r}K-\nabla^{2}\alpha+\alpha\left(A_{a}^{2}+2A_{b}^{2}+\frac{1}{3}\>K^{2}\right)$ (33)
$+$ $4\pi\alpha\left(\rho_{E}+S_{a}+2S_{b}\right),$
$\partial_{t}A_{a}$ $=$ $\beta^{r}\partial_{r}A_{a}-\left(\nabla^{r}\nabla_{r}\alpha-\frac{1}{3}\nabla^{2}\alpha\right)$ (34)
$+$ $\alpha\left({}^{(3)}R^{r}_{r}-\frac{1}{3}^{(3)}R\right)$
$+$ $\alpha KA_{a}-\frac{16}{3}\pi\alpha\left(S_{a}-S_{b}\right)\;,$
$\partial_{t}\hat{\Delta}^{r}$ $=$ $\beta^{r}\partial_{r}\hat{\Delta}^{r}-\hat{\Delta}^{r}\partial_{r}\beta^{r}+\frac{1}{a}\partial^{2}_{r}\beta^{r}+\frac{2}{b}\>\partial_{r}\left(\frac{\beta^{r}}{r}\right)$ (35)
- $\frac{2}{a}\left(A_{a}\partial_{r}\alpha+\alpha\partial_{r}A_{a}\right)$
$+$ $2\alpha\left(A_{a}\hat{\Delta}^{r}-\frac{2}{rb}\left(A_{a}-A_{b}\right)\right)$
$+$ $\frac{2\alpha}{a}\left[\partial_{r}A_{a}-\frac{2}{3}\>\partial_{r}K+6A_{a}\partial_{r}\chi\right$.
$+$ $\left.\left(A_{a}-A_{b}\right)\left(\frac{2}{r}+\>\frac{\partial_{r}b}{b}\right)-8\pi S_{r}\right]\;,$
where $\rho_{E}$ and $S_{r}$ are, respectively, the total energy density and the covariant component of the momentum density, as seen by an Eulerian observer, which are defined below. Note that the system of equations respects covariance, only that the terms of $\tilde{\gamma}^{rr}=a^{-1}$ and $\tilde{\gamma}^{\theta\theta}=\tilde{\gamma}^{\phi\phi}=1/(br^{2})$ have been explicitly written. The covariant derivatives with respect to the four-metric for the lapse function are computed as follows
$\nabla^{2}\alpha$ $=$ $\frac{1}{ae^{4\chi}}\left[\partial_{r}^{2}\alpha-\partial_{r}\alpha\left(\frac{\partial_{r}a}{2a}-\frac{\partial_{r}b}{b}-2\partial_{r}\chi-\frac{2}{r}\right)\right],$
$\nabla^{r}\nabla_{r}\alpha$ $=$ $\frac{1}{ae^{4\chi}}\left[\partial_{r}^{2}\alpha-\partial_{r}\alpha\left(\frac{\partial_{r}a}{2a}+2\partial_{r}\chi\right)\right]\;.$ (36)
$^{(3)}R^{r}_{r}$ and $^{(3)}R$ are the diagonal-radial component of the Ricci tensor and the Ricci scalar, respectively. Assuming spherical symmetry, they read
$^{(3)}R^{r}_{r}$= $-\frac{1}{ae^{4\chi}}\biggl[\frac{\partial^{2}_{r}a}{2a}-a\partial_{r}\hat{\Delta}^{r}-\frac{3}{4}\left(\frac{\partial_{r}a}{a}\right)^{2}+\frac{1}{2}\left(\frac{\partial_{r}b}{b}\right)^{2}$
$-\frac{1}{2}\hat{\Delta}^{r}\partial_{r}a+\frac{\partial_{r}a}{rb}+\frac{2}{r^{2}}\left(1-\frac{a}{b}\right)\left(1+\frac{r\partial_{r}b}{b}\right)$
$+4\partial^{2}_{r}\chi-2\partial_{r}\chi\left(\frac{\partial_{r}a}{a}-\frac{\partial_{r}b}{b}-\frac{2}{r}\right)\biggr],$ (37)
$^{(3)}R$= $-\frac{1}{ae^{4\chi}}\biggl[\frac{\partial^{2}_{r}a}{2a}+\frac{\partial^{2}_{r}b}{b}-a\partial_{r}\hat{\Delta}^{r}-\left(\frac{\partial_{r}a}{a}\right)^{2}$ (38)
$+\frac{1}{2}\left(\frac{\partial_{r}b}{b}\right)^{2}+\frac{2}{rb}\left(3-\frac{a}{b}\right)\partial_{r}b+\frac{4}{r^{2}}\left(1-\frac{a}{b}\right)$
$+8\left(\partial^{2}_{r}\chi+(\partial_{r}\chi)^{2}\right)-8\partial_{r}\chi\left(\frac{\partial_{r}a}{2a}-\frac{\partial_{r}b}{b}-\frac{2}{r}\right)\biggr].$
On the other hand, for the fluid dynamical system we use the framework described in Alcubierre’s textbook Alcubierre (2008), which uses the following covariant variables,
$D$ $\equiv\rho_{0}W~,$ (39)
$S_{r}$ $\equiv(\rho_{\rm fl}+p)W^{2}v_{r}~,$ (40)
$\mathcal{E}$ $\equiv(\rho_{\rm fl}+p)W^{2}-p-D~.$ (41)
Thus, is straightforward to identify the matter components in Eqs.(33-35) as
$\rho_{E}$ $=n^{\mu}n^{\nu}T_{\mu\nu}=D+\mathcal{E}~,$ (42)
$S_{i}$ $=-\gamma^{\mu}_{i}n^{\nu}T_{\mu\nu}=S_{r}~,$ (43)
$S_{ij}$ $=\gamma^{\mu}_{i}\gamma^{\nu}_{j}T_{\mu\nu}~,~,$ (44)
where from Eq.(44) we obtain
$S_{a}$ $\equiv\gamma^{rr}S_{rr}=\frac{1}{ae^{4\chi}}(E+D+p)+p~,$ (45)
$S_{b}$ $\equiv\gamma^{\theta}S_{\theta\theta}=r^{2}\frac{b}{a}~.$ (46)
The evolution equations for the fluid’s conserved variables {$E$, $D$, $S_{r}$}, written in terms of the BSSN variables above, read
$\left(\partial_{t}-\mathcal{L}_{\beta}\right)D$ $=-D_{k}(\alpha Dv^{k})+\alpha KD~,$ (47)
$\left(\partial_{t}-\mathcal{L}_{\beta}\right)S^{i}$ $=-D_{k}\left[\alpha\left(S^{i}v^{k}+\gamma^{ik}p\right)\right]$
$\>-(\mathcal{E}+D)D^{i}\alpha+\alpha KS^{i}~,$ (48)
$\left(\partial_{t}-\mathcal{L}_{\beta}\right)\mathcal{E}$ $=(\mathcal{E}+D+p)(\alpha v^{m}v^{n}K_{mn}-v^{m}\partial_{m}\alpha)$
$\>-D_{k}\left[\alpha v^{k}\left(\mathcal{E}+D\right)\right]+\alpha K(\mathcal{E}+p)~,$ (49)
where $D_{i}$ denotes the covariant derivative with respect to the spatial metric. A generic equation of state often requires a root-finding technique to obtain the value of $p$ and $v^{r}$, prior to the recovery of the other variables, as they are needed to evolve Eqs. (47)–(49). However, for cases with a simple equation of state, the recovery of $p$ and $v^{r}$ can be done analytically by finding the physical root of a high-order polynomial. For a barotropic fluid with $p=\omega\rho_{\rm fl}$, this implies solving a second-order polynomial Staelens et al. (2021), which diminishes the computational burden of the simulations.
A.2 Numerical Implementation
We numerically integrate the system of equations using a fourth-order Runge-Kutta (RK4) method. The spatial derivatives are discretized on a uniform grid using a standard five-point stencil, which provides fourth-order accuracy for the interior points. As customary, to ensure numerical stability in the presence of high-frequency modes, we apply Kreiss-Oliger dissipation.
Our computational domain is defined in logarithmic radial coordinates. We employ a grid of between $N=2000$ and $N=6000$ cells. The outer boundary of the grid is set to a comoving distance of $r_{\text{max}}\sim 20\,k_{*}^{-1}$, which corresponds to a physical size of approximately $2000\,H^{-1}(t_{0})$, ensuring it is sufficiently large to maintain a FLRW background at the outer boundary.
A.3 Construction of initial data
The initial data is computed by solving the Hamiltonian and Momentum constraints,
$\mathcal{H}$ $=~^{(3)}R+K^{2}-K_{ij}K^{ij}-16\pi\rho=0\,$, (50)
$\mathcal{M}_{i}$ $=D^{j}(K_{ij}-\gamma_{ij}K)-8\pi S_{i}=0\,$, (51)
where $^{(3)}R$ is the 3-Ricci scalar.
We assume a conformally flat initial hypersurface, corresponding to the BSSN variables $a=b=1$ and $A_{a}=A_{b}=\Delta^{r}=0$. Under these assumptions, the previous equations are solve by using
$K$ $=-\sqrt{24\pi\rho_{\rm bkg}(t_{0})}~,$ (52)
$\rho_{E}$ $=\rho_{\rm bkg}(t_{0})+\delta\rho~,$ (53)
$\delta\rho$ $=~^{(3)}R~,$ (54)
$S_{i}$ $=0~.$ (55)
We tested this choice of initial condition to another another popular choice based on gradient expansion Musco et al. (2005); Harada et al. (2015), and we obtained identical results for the thresholds and PBH masses. However, because our method provides an exact solution to the constraint equations, rather than introducing deviations of order $\mathcal{O}(\epsilon^{2})$ from the gradient expansion, the constraints are better satisfied throughout the numerical evolution.